OpenStudy (anonymous):

A community is building a square park with sides that measure 150 meters. To separate the picnic area from the play area, the park is split by a diagonal line from opposite corners. Determine the approximate length of the diagonal line that splits the square. If necessary, round your answer to the nearest meter.

OpenStudy (anonymous):

Options: 150 meters 300 meters 212 meters 45,000 meters

OpenStudy (anonymous):



OpenStudy (anonymous):

do you think a 300 meter line would fit there?

OpenStudy (anonymous):

See the diagonal of the square is given by: \[d = \sqrt{2}side\] SO: \[d = \sqrt{2}\times 150 = 212.132 =>212\] @babylailuv

OpenStudy (anonymous):

@babylailaluv